OpenStudy (anonymous):

(-8-11i)+(-1+i) -4i(3-i) (2-11i)-(10+6i) -i(4i) (3+8i)-(-2-i) (3+5i)(2-4i) PLEASE HELP:)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

HELP SOMEONE:)plz

OpenStudy (zzr0ck3r):

do the multiplication and then add them up, i * i = -1

OpenStudy (zzr0ck3r):

ill do the -i*4i*(3+8i) -i*4i*(3+8i) = -*-4(3+8i) = 12+24i

OpenStudy (anonymous):

wow thank you:P